OpenStudy (anonymous):

Use the limit definition of the derivative. f(x)=1/x+3 f'(6)=?

OpenStudy (anonymous):

\[f'(x)=\lim_{x\to 6}\frac{\frac{1}{x+3}-\frac{1}{9}}{x-6}\]

OpenStudy (anonymous):

A derivative of a function f(x) is defined as the limit of a ratio of two expressions as x approaches a value c, in your case c=6.

OpenStudy (anonymous):

then a bunch of algebra \[\frac{1}{x+3}-\frac{1}{9}=\frac{9-(x+3)}{9(x+3)}=\frac{6-x}{9(x+3)}\]

OpenStudy (anonymous):

divide by \(x-6\) to give you \[-\frac{1}{9(x+3)}\] and finally replace \(x\) by \(6\) for your answer

OpenStudy (anonymous):

I think a general form of the definition of a derivative should be helpful.

OpenStudy (anonymous):

I am confused. i dont get the first line of the equation why its( 1/x+3-1/9)/x-6?

OpenStudy (anonymous):

and the answer is -(1/81)?